Q9 : A car covers 300 m in 20 seconds, while accelerating
uniformly at a rate of 5m/s2. Determine :
(i) Initial and final velocities of the car
(ii) Distance travelled before coming to this point assuming it started
from rest
(iii) Its velocity after the next 10 seconds. Q10 : A tennis ball traveling at 25.0 m/s is returned by Williams. It
leaves his racket with a speed of 39.0 m/s in the opposite direction from
which it came. i) Calculate the change in momentum of the tennis ball.
ii) If the 0.080 kg ball is in contact with the racket for 0.020 s,
calculate the average force needed for Williams to hit the ball.
iii) As a Safety Officer predict what happens if the mass of the ball
increases.
1
Expert's answer
2019-12-13T09:52:55-0500

9.

(i)

"v^2-u^2=2as=2\times5\times300=3000\ ......(1)"

also,

"v-u=at=5\times20=100\ ...(2)"

solving (1) and (2)

We get

"v=65\ m/sec" and "u=-35\ m/sec"

(ii)

"s=\frac{1}{2}at^2=0.5\times5\times20^2=1000\ m"

(iii)

"v=u+at=-35+5\times10=15\ m/sec"


10.

(i)

change in momentum = "m(v_2-v_1)=0.080\times(-39-25)=-5.12\ N\ m/sec"

(ii)

average force needed = "\frac{\Delta P}{\Delta t}=\frac{5.12}{0.020}=256\ N"


(iii)

If the mass of ball increses then william will have to apply more force to return it due to which he can damage his shoulder.